Solve the system by addition.
3x – y = 1
3x – y = 2

Answer by checkley71(8403)

3X-Y=1 NOW MULTIPLY THE OTHER EQUATION BY -1 & ADD.
-(3X-Y=2)
--------------------------------------
0X+0Y=-1
0=-1 NO SOLUTION BY ADDITION.
--------------------------------------